OpenStudy (anonymous):

How do you work 7q+q-3q=-24

OpenStudy (mathstudent55):

Add the like terms on the left side.

OpenStudy (anonymous):

Then divide to isolate the q.

OpenStudy (anonymous):

Just as 7+1-3=5, 7q+q-3q=5q. So... 5q = -24, and q = \[\frac{ -24 }{ 5 }\]

OpenStudy (mathstudent55):

Like terms are terms that have the same variable part.

OpenStudy (anonymous):

Do you get it?